A stock had returns of 6 percent, -22 percent, 18percent, 12 percent, and -2 percent over the past five years. What is the standard deviation of these returns?
Answer #1

Average return=Total return/Total time period

=(6-22+18+12-2)/5=2.4%

Return (Return-Average Return)^2
6 (6-2.4)^2=12.96
-22 (-22-2.4)^2=595.36
18 (18-2.4)^2=243.36
12 (12-2.4)^2=92.16
-2 (-2-2.4)^2=19.36
Total=963.2%

Standard deviation=[Total (Return-Average Return)^2/(Time period-1)]^(1/2)

=[963.2/(5-1)]^(1/2)

=[963.2/4]^(1/2)

=15.52%(Approx).